A fire truck is emitting a 450 Hz signal passes by a stationary detector. The difference in frequency measured by the detector is 58 Hz. If the speed of sound is 345 m/s, how fast is the truck moving? (ans: 22m/s)


1
Expert's answer
2022-05-18T08:31:42-0400

The Doppler's effect says


"f=f_0\frac{c\pm v_r}{c\pm v_s}""f=f_0\frac{c}{c+ v_s}""v_s=c(f_0/f-1)\\
=345(450/(450-58/2)-1)=24\:\rm m/s"